I need to figure out how to put this function y=x^2-4x-5 into this form
y=a(x-h)^2+k

Answer by dolly(163)

Given y = x^2 - 4x - 5
==> y = x^2 - 4x + 4 - 4 - 5 [add and subtract 4 to make the first3 terms a perfect square]
==> y = (x^2 - 4x + 4) - 9
==> y = (x - 2)^2 - 9
This is of the form y = a(x-h)^2 + k where
a = 1; h = 2 and k = -9